OpenStudy (anonymous):

1.A plane travels from Orlando to Denver and back again. On the five-hour trip from Orlando to Denver, the plane has a tailwind of 40 miles per hour. On the return trip from Denver to Orlando, the plane faces a headwind of 40 miles per hour. This trip takes six hours. What is the speed of the airplane in still air?

OpenStudy (anonymous):

i know the equation to solve it is 5 * (a+40)= 6 * (a-40) but idk how to solve it step by step

OpenStudy (anonymous):

if this eq. is correct then use the property a(b+c)=ab+ac.

OpenStudy (anonymous):

and how do i do that? /:

OpenStudy (anonymous):

from the above property we can write it as 5*a + 5*40 = 6*a - 6*40

OpenStudy (anonymous):

= 5a+200=6a-240 @itsmichi95 understood?